# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

### v2

@@ 4. Background
Table 2. Average F-scores and percentage of rejected null hypotheses $H_{0}$ ($p=0.05$) across different lag values, comparing the standard datasets with the ODE benchmark. Bold values indicate the highest correlation at each lag.
Dataset Type (lag) Average F-score $H_{0}$ rejected (%, p=0.05)
@@
Table 3. Average MSE across four forecasting horizons for each model variant on the ODE Benchmark. Bold values highlight the best variant within each model; the best overall result per row is shown in blue.
@@
Fixing Lookback Windows:
The problem formulation described earlier places no intrinsic requirement on the forecasting query $\bm{X}$} to have a fixed lookback length $L$. Nevertheless, most recent LTSF models fix the lookback window during training, typically to a value of $L=96$. This constraint is usually imposed for implementation convenience, computational efficiency, or alignment with benchmarking conventions, rather than for a principled reason. As we have argued, the lookback window should instead be treated as an integral property of the model, similar to any other model-specific hyperparameter.
Channel Independence:
One of the main findings of PatchTST (Nie et al., 2023) and later by TSMixer (Chen et al., 2023) is that, on widely used benchmark datasets, CI models are sufficient for accurate forecasting. This suggests that interactions between different channels are not essential, and that each channel can be modeled in isolation. Formally, a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times H}$ is said to be channel-independent if for all ${\bm{X}}\in\mathbb{R}^{C\times L}$ and any channel input ${X}_{c}\in\mathbb{R}^{L}$,
(2) $\hat{{\bm{Y}}}({\bm{X}})_{c}=\hat{{\bm{Y}}}({{X}}_{c})$
i.e., the prediction for any given channel $c$ is generated using only its own historical data, independent of all other channels.
Granger Causality:
The Granger causality test (Granger, 1969) evaluates whether past values of $X_{c_{2}}$ improve forecasts of $Y_{c_{1}}$ beyond what is achievable using only past values of $X_{c_{1}}$. If including $X_{c_{2}}$ reduces prediction error in an autoregressive model, then $X_{c_{2}}$ provides a useful linear signal for forecasting channel $c_{1}$. To quantify this directional relationship, we compute the F-statistic (Equation˜5) for each channel pair, measuring the predictive gain from adding $X_{c_{2}}$.
The F-statistic compares the sum of squared residuals (SSR) of two models: a univariate model using only lagged values of $c_{1}$ ($SSR_{u}$, Equation˜3), and a multivariate model that additionally includes lagged values of $c_{2}$ ($SSR_{mv}$, Equation˜4). The univariate model uses coefficients $\hat{a}$ applied to $X_{c_{1}}^{t-i}$ over a lag window of length $L$, where $i=1,\dots,L$; the multivariate model extends this with coefficients $\hat{b}$ for $X_{c_{2}}^{t-i}$. Residuals are computed for $t=L+1,\dots,L+H$.
In Equation˜5, $k_{u}$ is the number of added parameters associated with channel $c_{2}$, and $k_{mv}$ is the total number of parameters in the multivariate model. The denominator degrees of freedom are $N-k_{mv}$, where $N$ is the number of observations. A large F-statistic indicates a significant improvement in explained variance when including $c_{2}$. We reject the null hypothesis $H_{0}\!:$$c_{2}$ does not Granger-cause $c_{1}$” when this improvement is statistically significant.
(3) $SSR_{u}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}\bigg)^{2},$
(4) $SSR_{mv}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}-\sum_{i=1}^{L}\hat{b}_{i}X_{c_{2}}^{t-i}\bigg)^{2},$
(5) $F=\frac{(SSR_{u}-SSR_{mv})/k_{u}}{SSR_{mv}/(N-k_{mv})},$
5. General Experimental Setup
5.1. Models
We evaluate a diverse set of strong baseline models, varying in architectural complexity and treatment of channel dependencies. These include MLP-based models (TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024b)), attention-based models (PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024), Crossformer (Zhang and Yan, 2023)), and the simple yet competitive DLinear (Zeng et al., 2023). We use only the CI version of DLinear, consistent with its design as a lightweight baseline. For iTransformer, we use only the CD version, as removing inter-channel attention would reduce it to a basic MLP, diverging from its core design. Rest of the methods are available in or allow derivation of both CD and CI forms. Additional details on the derivation of both variants are included in the appendix. For additional experimental details, please refer to our code repository [^1] https://github.com/anonsub936/How-Biased-is-Time-Series-Forecasting-.
5.2. Datasets
The Standard Datasets:
The first category includes widely used benchmarks from recent state-of-the-art studies and surveys (Nie et al., 2023; Liu et al., 2024; Wang et al., 2024b; Chen et al., 2023; Zhang and Yan, 2023; Zeng et al., 2023): Weather, Electricity, Traffic, and the Electricity Transformer Temperature (ETT) datasets. These benchmarks gained popularity following their adoption in the Informer paper (Zhou et al., 2021). We refer to this group as the standard datasets. Details on data statistics are provided in the appendix.
The Chaotic ODE Benchmark:
As a second dataset collection, we consider the benchmark proposed by Gilpin (2021). This benchmark was introduced as a time-series collection that features complex time series dynamics with clear real world data properties, providing a challenging and realistic setting for channel-dependent time-series forecasting. A demonstration for one of these datasets and how it simulates a recurrent patterns across different channels can be seen in Figure˜2. This figure particularly emphasizes periods of synchronization in the angular velocities of the two pendulums during the final 200 timesteps, indicative of inter-channel correlation. Additional plots of the selected ODE system are provided in the appendix.
Figure 2. Time series visualization of the Double Pendulum ODE system. Notably, the bottom two channels (angular velocities of the two pendulums) exhibit synchronized behavior in the last 200 time steps, demonstrating time-varying inter-channel correlation.
We select a subset of datasets from this benchmark based on two distinctive criteria: (i) The datasets follow complex trajectories that challenge most time series models and (ii) they are derived directly from the underlying ODEs, ensuring a clear mathematical correlation between variables. For criterion (i), we use the Largest Lyapunov exponent (Gilpin, 2021) to quantify dataset complexity, selecting datasets with high values to ensure forecasting remains non-trivial task that cannot be solved easily based on one channel information. To illustrate criterion (ii), we show below a simplified form of the Double Pendulum system (Elbori and Abdalsmd, 2017), assuming equal masses and rod lengths. The coupling between angular velocities $\frac{d\theta_{1}(t)}{dt},\frac{d\theta_{2}(t)}{dt}$ and angles $\theta_{1}(t),\theta_{2}(t)$ at time $t$ from Equations˜6 and 7 is evident, with parameters $g$ (gravity) and $l$ (rod length) incorporated. Such systems are essential in modeling real-world physics applications, including robotics (Jadlovská and Sarnovskỳ, 2012).
Due to space constraints, we defer detailed discussion of another example, Cell Cycle, to the appendix. This dataset reflects strong variable coupling and captures real-world measurements of Cyclin-Dependent Kinase (CDK) activity, crucial to cell cycle progression (Romond et al., 1999). Our selection yields six datasets with rich temporal dynamics and strong inter-channel dependencies, referred to as the ODE datasets. Further reasoning for particularly choosing this benchmark follows through the statistical analysis in Section˜6.2.
(6) $\frac{d^{2}\theta_{1}(t)}{dt^{2}}$ $=-\frac{2g}{l}\theta_{1}(t)+\frac{g}{l}\theta_{2}(t)$
(7) $\frac{d^{2}\theta_{2}(t)}{dt^{2}}$ $=\frac{2g}{l}\theta_{1}(t)-\frac{2g}{l}\theta_{2}(t)$
5.3. Hyperparameter Tuning
To ensure fair evaluation, we apply a consistent tuning procedure across all models. Each hyperparameter is defined either by a discrete set (categorical) or a continuous range (real-valued). Performance for each hyperparameter configuration is validated on the designated validation set through the Mean Squared Error (MSE). The best model is chosen based on the lowest validation error, then it is tested on a the separate test split to get the final evaluation. We use Bayesian optimization via Optuna (Akiba et al., 2019) to efficiently explore promising configurations, running 20 trials per setup (i.e., a specific model, dataset, and forecasting horizon).
Crucially, the lookback window is treated as a tunable hyperparameter, selected from $\{96,192,336,512,720\}$. Also more hyperparameters that are specific for each model are tuned thoroughly, allowing each model to operate under optimal conditions for fair comparison. Detailed search spaces and best hyperparameters chosen for each model are provided in the appendix.
5.4. Granger Causality
We conduct Granger causality analysis following the standard procedure outlined by Granger (1969), with several preprocessing steps to ensure robust and meaningful results:
  1. (i)
    We first assess the stationarity of each dataset. If non-stationarity is detected, differencing is applied iteratively until stationarity is achieved.
  2. (ii)
    To remove redundancy, we compute the Pearson correlation coefficient for all channel pairs. Channels with correlation exceeding 0.95 are considered highly similar, and one from each such pair is excluded from further analysis.
After the preprocessing steps, we conduct the F-test described in Section˜4. Consequently, we report two important metrics: the average F-score over all valid channel pairs, and the percentage of pairs for which the null hypothesis is rejected (indicating significant Granger causality). These results are summarized in Section˜6.2. Additional implementation details, including parameter settings and the formal hypothesis definition, are provided in the appendix.
6. Results
6.1. Evaluation of Standard Datasets
We reproduce the current time-series forecasting models in our above-mentioned evaluation settings, where we conduct fair and extensive hyperparameter tuning, including tuning the lookback window. Unless stated otherwise, all results are averaged over 5 different random seeds. Table˜1 presents the results averaged across forecasting horizons 96, 192, 336, and 720 for the standard datasets. Detailed values for each individual horizon and standard-deviations are provided in the appendix.
PatchTST remains the state-of-the-art for time-series forecasting. In 3 out of 7 datasets, the standard channel-independent version of PatchTST achieves the best performance. For three other datasets, this version ranks second or third. Its performance is not among the top three only on the Weather dataset, which may be attributed to the relatively stronger inter-channel dependencies in this dataset, as indicated by Granger causality. The average rank of PatchTST across all datasets is 2.14 showing a significant gap of 0.57 compared to the second-best model, the channel-independent version of TSMixer.
Linear models, meanwhile, remain competitive with recent forecasting approaches. The performance of DLinear on the standard datasets is comparable to more recent models. Across the 6 datasets, it achieves an average rank of 5.43 out of 10. Notably, on ETTh1, it ranks second, trailing PatchTST by only a small margin.
6.2. Granger Causality for Evaluation of Channel Correlations
Based on the findings from the previous section, which showed that multivariate models do not consistently outperform univariate ones on the standard datasets, we turn to a statistical evaluation of channel correlations across a broader set of datasets. Our goal is to identify benchmarks where inter-channel dependencies exist, potentially finding datasets and evaluation setup where multivariate models are needed.
Although the full test results are presented in the appendix, we focus here on a representative comparison between the ODE datasets and the standard datasets, to highlight key patterns. We conduct Granger causality tests using three different lag values: 30, 96, and 192. These lags correspond to different lookback window sizes used in the autoregressive models underlying the tests. To further analyze channel dependence across a broad range of time series domains, we computed Granger causality values for 41 real-world datasets from the extensive Monash Time Series Forecasting Archive (Godahewa et al., 2021). These results, calculated for lag values of 7, 30, 96, and 192, are provided in the appendix. Throughout this section, we use the terms lag and lookback window interchangeably.
Varying the lag allows us to analyze how inter-channel correlations evolve with different temporal resolutions. We can see the F-scores and the percentage of pairs of channels that exhibits granger causality in Table˜2. For both metrics, the higher values indicate higher correlation in the datasets. As seen in the table, the ODE datasets exhibit significantly higher F-scores and pass rates across all lag values, strongly indicating the presence of meaningful inter-channel dependencies. On contrary, the standard datasets show considerably weaker correlations, with average F-scores dropping below 1.0 for ETTm2 dataset at the highest lag.
These findings align with our earlier empirical observations: univariate models often match or outperform multivariate models on the standard datasets due to weak inter-channel coupling. Moreover, we observe that increasing the lag generally reduces the measured correlation for both dataset types. This effect is particularly pronounced for the standard datasets, suggesting that longer lookback windows benefit more univariate models than multivariate ones, since the former cannot leverage strong inter-channel signals.
In contrast, the ODE datasets show much stronger correlations at longer lags, e.g., average F-scores of 88.5 (lag 96) and 35.29 (lag 192), compared to 1.40 and 1.23 for the standard datasets. These results support using the ODE datasets as a valuable testbed for evaluating models under strong cross-channel dynamics.
6.3. Evaluation on the ODE Benchmark
As discussed in Section˜6.1, our results reaffirm DLinear as a competitive baseline and also suggest the superiority of CI models over their CD counterparts. In this section, we evaluate how the strong correlations between channels in the ODE datasets affect model performance. Table˜3 presents the results obtained using the same evaluation protocol described in Section˜6.1.
The evaluation on the ODE datasets reveals a markedly different performance hierarchy among models. First of all, CD variants clearly outperform CI ones. The CD variants of the evaluated models consistently outperform their CI counterparts, highlighting the importance of modeling inter-channel dependencies, especially in datasets governed by coupled differential equations. As shown in Table˜4, CD models outperform CI versions in 18 out of 24 total experiments. This strongly supports the claim that model architecture should be tailored to the underlying data structure. Secondly, Linear models perform poorly on the ODE benchmark. Contrary to their performance on standard datasets, linear models such as DLinear are no longer competitive here. Averaged over all forecasting horizons, DLinear ranks worst in 5 out of the 6 ODE datasets, with an average rank of 9.83. Compared to Crossformer, DLinear suffers from performance drops ranging from 45% to 118%, with a median degradation of 58.3% and a mean of 74%.
In summary, the commonly held view that linear models are competitive with deep architectures does not generalize to datasets with strong inter-channel dynamics. Our findings underline the importance of evaluating forecasting models on datasets with diverse structural properties. In cases where inter-channel correlation is intrinsic to the data—as with ODE systems—model architectures must be explicitly designed to capture these dependencies in order to perform well.
6.4. Lookback Window Analysis
In this section, we extend our evaluation by analyzing the distribution of selected lookback window values, grouped by both datasets and models. This analysis highlights how an optimal lookback window varies significantly depending on the model type and dataset characteristics.
As shown in Figure˜3, we report the frequency of best-performing lookback window across CI models, CD models, and datasets. In the Figure 3(a), CI models exhibit substantial variation in their preferred lookback windows. For instance, DLinear consistently favors longer windows (336 or 720), while shorter windows are rarely optimal. This finding is soldified by the fact that overall, CI models choose 96 lookback window only 11% times. In contrast, Figure 3(b) shows that CD Models more often benefit from shorter windows, with 96 or 192 selected in 55% of the cases. These counts are aggregated across the four forecasting horizons (96, 192, 336, and 720).
Figure 3(c) shows that optimal lookback windows also vary widely by dataset. This aligns with expectations: for hourly datasets like ETTh1 and ETTh2, a window of 96 covers four days, often sufficient for effective forecasting. In contrast, for 15-minute interval datasets like ETTm1 and ETTm2, more timesteps are required to capture similar temporal spans. Accordingly, ETTm2 frequently favors longer windows larger than or equal to 336, while ETTh2 consistently prefers windows shorter than 336.
Overall, this analysis reinforces the importance of tuning the lookback window, not just across model types, but also in response to dataset-specific temporal characteristics. For per-horizon results, refer to the bar charts in the appendix.
@@ 6.4. Lookback Window Analysis
Figure 3. Frequencies of best-performing lookback windows grouped by: (a) CI models, (b) CD models, and (c) datasets.
7. Conclusion and Recommendations
7.1. Conclusion
Model ODE-6 Standard-7
CI CD CI CD
PatchTST 4 2 6 1
TSMixer 0 6 7 0
Crossformer 0 6 4 3
TimeMixer 2 4 4 3
Totals 6 18 21 7
Table 4. Frequency of CI and CD variants outperforming each other across models and datasets.
In this paper, as outlined in the introduction, our primary goal is to establish a fair and unbiased evaluation framework for multivariate TSF models. To achieve this, we make several critical adjustments to the experimental setup including: tuning the lookback window, analyzing statistical inter-channel correlations, and conducting empirical evaluations on datasets with varying characteristics. From this comprehensive setup, we draw the following key conclusions:
  1. (1)
    Tuning the lookback window is essential for fairness in multivariate TSF. Fixing the lookback window (commonly at 96) can unfairly disadvantage univariate models. Allowing models to tune this parameter ensures a more balanced comparison and reveals meaningful performance differences. This insight directly enables the second key finding below.
  2. (2)
    Modeling channel dependencies is unnecessary for standard datasets. For all models where a CI variant is available, it consistently outperforms the CD counterpart, by a margin of 21 to 7 on standard benchmarks. This is a significant result, especially considering that models like TSMixer, Crossformer, and TimeMixer are inherently designed to mix information across channels. Our findings indicate that inter-channel mixing is not a major contributor to performance on these datasets and can, in fact, degrade it (Table˜4).
  3. (3)
    Dataset nature plays a critical role in multivariate TSF. This aspect has been largely overlooked in prior work. Our experiments (Table˜4) reveal that datasets with different underlying dynamics require different modeling strategies. For instance, ODE Datasets exhibit significantly higher inter-channel correlation, as confirmed by statistical tests. This justifies the use of multivariate models in such cases and is further supported by the superior performance of channel-dependent models on this dataset family.
7.2. Recommendations for Multivariate Time-Series Forecasting Evaluation
In this paper, we strive to come with a useful insights that can advance the field of multivariate time series forecasting and this was done through a detailed experiments and statistical analysis from which we draw the following recommendations:
  • Do not use a fixed lookback window when a vast amount of historical data is available to ensure full potential of all models considered as seen in Section˜6.1 and Section˜6.3.
  • For proper evaluation and before choosing model architecture, consider statistically testing the channel correlation of the datasets as shown in Section˜6.2, it can help a lot in making an informed decision between deploying a CD or a CI model.
  • As noted in Section˜6.4, shorter lookback window favors in general CD models, which means for applications with limited amount of historical data, an architecture capturing channel dependency is generally recommended.
Appendix A Description of All Datasets
In this section, we would like to show in detail the different datasets used in our experiments. Starting with the standard datasets which were explained a lot in literature, so we will just give a brief description of them. First ETT- datasets which representing indication of long term power deployment which are recorded for 2 different granalurities (15 mins for ’m’ and 1 hour for ’h’) for 2 different counties (hence the post-fix ’1’ and ’2’). Electricity dataset follows the consumption of electricity over an extended amount of time (3 years) recorded hourly for 321 different households. Traffic dataset is representing the flow of traffic recorded through multiple sensors in an hourly frequency spanning over a period of approximately 2 years. All of this information as well as the train/validation/test split is shown in Table˜5
Dataset/Feature Channels Granularity Train/val/test Total Timesteps
ETTh (1/2) 7 1 hour 12/4/4 (months) 17420
ETTm (1/2) 7 15 mins 12/4/4 (months) 69680
electricity 321 1 hour 7:1:2 26304
traffic 862 1 hour 7:1:2 17544
weather 21 10 mins 7:1:2 52696
Table 5. Basic statistics about the standard datasets widely used in the time series forecasting literature. This table shows the number of features, frequency of recording the data, train/validation/test split as well as the total timesteps recorded in a given dataset.

### v3

@@ 6.1 Evaluation of Standard Datasets
6.2 Granger Causality for Evaluation of Channel Correlations
Building on previous findings that multivariate models do not consistently outperform univariate models on standard datasets, we perform a statistical analysis of channel correlations across both standard datasets and ODE datasets to determine where inter-channel dependencies exist and multivariate models might be necessary.
We compare both benchmarks using Granger causality tests with lags (lookback windows) of 30, 96, and 192. F-scores and percentages of channel pairs showing causality (i.e. $H_{0}$ rejected) are reported in Table˜3, with higher values indicating stronger correlations. The ODE datasets show significantly stronger inter-channel dependencies with higher F-scores and $H_{0}$ rejection at all lags, while standard datasets have much weaker correlations, with some F-scores dropping below 1.0 at higher lags. This supports and explains earlier observations where univariate models often match or outperform multivariate ones on standard datasets due to weak coupling.
Increasing lag generally reduces correlations in both dataset types, especially for standard datasets, suggesting longer lookback windows favor univariate models that do not use cross-channel signals. In contrast, ODE datasets maintain strong correlations even at longer lags (F-scores of 88.5 at lag 96 and 35.29 at lag 192), supporting their use as testbeds for models handling strong cross-channel dynamics.
Table 3: Average F-scores and $H_{0}$ rejection rates ($p=0.05$). Bold indicates the highest correlation at each lag.
Dataset Type (lag) Average F-score $H_{0}$ rejected (%)
@@ 6.2 Granger Causality for Evaluation of Channel Correlations
6.3 Evaluation on the ODE Benchmark
As discussed in Section˜6.1, DLinear remains a competitive baseline on standard datasets, with CI models outperforming CD counterparts. However, on the ODE datasets, model performance differs significantly. CD variants consistently outperform CI models in 18 out of 24 experiments as shown in Table˜4, underscoring the importance of modeling inter-channel dependencies in datasets governed by channel-interactions.
Interestingly, Linear models perform poorly on ODE datasets. DLinear ranks worst in 5 of 6 ODE datasets, with an average rank of 9.83. Compared to Crossformer, DLinear’s performance drops between 45% and 118%, with a median decline of 58.3%.
These results challenge the notion that linear models compete well with deep architectures, highlighting the need to evaluate models on datasets with varying structural characteristics. For data with intrinsic inter-channel correlations, like ODE systems, model architectures must explicitly account for these dependencies to achieve strong performance. We note that while Granger causality tests provide a rigorous measure of linear statistical predictability between channels, they serve as a proxy for inter-channel dependency in this context rather than implying structural or physical causality in the strong sense.
Table 4: Mean MSE on the ODE Benchmark. Bold: best per model; blue: overall best. Superscript references statistical details in the extended version.
@@ 6.3 Evaluation on the ODE Benchmark
6.4 Lookback Window Analysis
This section analyzes the distribution of optimal lookback window values by dataset and model, showing significant variation. Figure˜3 shows that CI models favor longer windows (336 or 720), with 96 chosen only 11% of the time. In contrast, CD models often benefit from shorter windows (96 or 192) in 55% of cases.
Theoretically, this divergence suggests a trade-off between temporal depth and inter-channel noise. CI models, treating each channel in isolation, require longer historical contexts to compensate for the lack of cross-variable information. Conversely, CD models extract relevant signals from other channels, allowing them to achieve high accuracy with shorter temporal windows. Furthermore, for CD models, excessively long windows may introduce redundant inter-channel noise, making it harder for the model to distinguish true dependencies from spurious correlations across variables.
Comparing across datasets, hourly datasets (ETTh1 and ETTh2) typically prefer shorter windows around 96, covering approximately four days, while 15-minute interval datasets (ETTm1 and ETTm2) favor longer windows (greater than 336) to capture equivalent temporal spans. This highlights that while temporal span is a physical requirement of the data, the optimal window length is a function of the model’s capacity to process spatial (channel) versus temporal information.
@@ 6.4 Lookback Window Analysis
Figure 3: Frequencies of best-performing lookback windows grouped by: (a) CI models, (b) CD models, and (c) datasets.
7 Conclusion and Recommendations
Our study establishes a fair evaluation framework for multivariate time series forecasting by tuning lookback windows, analyzing inter-channel correlations, and assessing diverse datasets. From this, we conclude and recommend the following:
  1. 1.
    Lookback window tuning is essential. Fixing the lookback window (commonly at 96) biases results against univariate models and masks true performance differences. Recommendation: Always allow models to optimize lookback windows to ensure balanced and meaningful comparisons.
  2. 2.
    CD modeling is unnecessary for standard datasets. CI models generally outperform CD models on standard datasets by a margin of 21 to 7, suggesting that mixing information across channels can negatively impact performance. Recommendation: For standard datasets, CI models should be preferred as they typically provide better results.
  3. 3.
    Dataset characteristics dictate model design. Datasets with strong inter-channel correlations, such as ODE datasets, require CD models capturing these dependencies, supported by superior CD performance and statistical tests. Recommendation: Pre-evaluate channel correlations using statistical tests to select appropriate model architectures and use CD models on datasets exhibiting strong inter-channel dynamics.
  4. 4.
    Lookback window length interacts with channel dependency. CD models benefit from shorter lookback windows, while CI models may favor longer ones, reflecting differing temporal dependency needs. Recommendation: Tailor lookback window length jointly with model type and dataset temporal resolution to maximize forecasting accuracy.
Limitations and Future Work:
While our framework ensures a fair comparison, it is subject to several limitations. First, our evaluation relies on a predefined discrete set of lookback windows $\{96,192,336,512,720\}$; while this aligns with standard practice in LTSF literature, it remains a pragmatic subset of all possible temporal spans. Second, our primary reliance on MSE, while standard for penalizing large outliers, may not fully capture model performance in contexts where distributional or scale-invariant metrics are prioritized.
Furthermore, while this study leverages chaotic ODE simulations to provide a mathematically verifiable ground truth for channel dependency, future work will explore whether these advantages transfer to highly stochastic real-world domains—such as high-frequency financial markets or complex human behavior analysis—to further validate the broader applicability of our findings. Finally, our observations regarding CI and CD model performance are conditioned on the current state-of-the-art model families; the emergence of hybrid architectures may shift these dynamics in future research.
Appendix 0.A Description of All Datasets
In this section, we provide a detailed description of the various datasets used in our experiments.
Standard Datasets
We first utilize standard benchmarks widely recognized in the time series forecasting literature. The ETT (Electricity Transformer Training) datasets represent long-term power deployment indicators, recorded at two granularities: 15-minute intervals (denoted by ’m’) and 1-hour intervals (denoted by ’h’) across two different counties (1 and 2). The Electricity dataset tracks hourly consumption for 321 households over three years. The Traffic dataset represents traffic flow recorded by 862 sensors at an hourly frequency over two years. Finally, the Weather dataset includes 21 meteorological features recorded every 10 minutes. Statistics for these datasets are summarized in Table 5.
Table 5: Basic statistics about the standard datasets used in the time series forecasting literature.
Dataset Channels Granularity Split (T/V/T) Total Steps
ETTh (1/2) 7 1 hour 12/4/4 (months) 17,420
ETTm (1/2) 7 15 mins 12/4/4 (months) 69,680
Electricity 321 1 hour 7:1:2 26,304
Traffic 862 1 hour 7:1:2 17,544
Weather 21 10 mins 7:1:2 52,696
Chaotic ODE Benchmark